OpenStudy (agentnao):

I have another derivative question, this time involving the chain rule. Can anyone help me?

OpenStudy (anonymous):

?

OpenStudy (agentnao):

\[(\sin x/1+\cos x)^{2}\] They want me to solve this using the chain rule, but I'm not really sure how to go about it.

OpenStudy (anonymous):

\[\left(\frac{\sin(x)}{1+\cos(x)}\right)^2\]?

OpenStudy (agentnao):

Yes! That's the question.

OpenStudy (anonymous):

you are going to take the derivative of the square first, the derivative of the squaring function is the doubling function, so the first thing you write is \[2\left(\frac{\sin(x)}{1+\cos(x)}\right)\times \frac{d}{dx}\left(\frac{\sin(x)}{1+\cos(x)}\right)\] but you still have to take the derivative of that quotient, and there is no way to do with without the quotient rule

OpenStudy (anonymous):

you need help with that part or is that one ok?

hartnn (hartnn):

there is a way to do it without quotient rule, but for that u need to be strong in trigonometry....

OpenStudy (agentnao):

I understand the first part, so now for the second part, the quotient rule is really all I have to do?

OpenStudy (anonymous):

yes

OpenStudy (agentnao):

Let me work it out, then I'll send you my answer and you can tell me if it's right!

OpenStudy (anonymous):

i mean "all" is not how i would put it because the quotient rule is kind of a drag, but yes, that is what you have to do

OpenStudy (agentnao):

Haha, definitely a drag. I think I've got it though. Is the answer going to be pretty long for this one?

OpenStudy (anonymous):

well not really there is a raft of cancellation in the numerator i think just grind it out, it is not that bad

OpenStudy (agentnao):

OH!! Oh my goodness, I see what you mean. Is this it? \[\frac{ 2\sin (x) }{ (1+\cos (x))^{2} }\]

OpenStudy (anonymous):

yes

OpenStudy (agentnao):

Thank you so much! I'll probably be asking another similar question shortly, but thank you for helping me with this one :)

OpenStudy (anonymous):

your welcome, hope the hurricane passed you by

OpenStudy (agentnao):

It did, thank goodness! I'm completely down south, so it didn't affect me at all, but I feel so terrible for everyone up north! I hope you weren't affected either!

OpenStudy (anonymous):

power went out for like 2 second, but all around was chaos, just not right where i live same thing happened with irene just lucky twice i guess

OpenStudy (agentnao):

\[5\cot \(\frac{ 2 }{ x })\] Is the answer to ^^ that \[10x ^{-2}\csc ^{2}(\frac{ 2 }{ x })\] ?

OpenStudy (agentnao):

Wow! Lucky indeed! I hope everyone else can make a speedy recovery, but the damage looked pretty extensive :/

hartnn (hartnn):

thats correct.

OpenStudy (agentnao):

Sweet! Thank you :)